40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 12 20:19:25 2016, Lydia, Update, General, MC REFL beam splitter not replaced 
    Reply  Wed Sep 14 19:41:23 2016, Johannes, Update, General, PSL back to high power 
Message ID: 12493     Entry time: Wed Sep 14 19:41:23 2016     In reply to: 12485
Author: Johannes 
Type: Update 
Category: General 
Subject: PSL back to high power 

Today's summary:

  • Replaced mirror in MC REFL path with R=10% BS and aligned beam on PD while still at low power
  • Replaced HR mirror in Transmon path at EY table with 50/50 BS. Alignment onto QPD not yet confirmed because we need IR from the YARM for it.
  • Put ND filters back on Transmon QPDs at both X and Y ends. For now I put all the filters on, for a combined OD of 1.6 at both ends (1.0 + 0.6 at YEND and 1.0 + 0.4 + 0.2 at XEND).
  • Put ND filter back on Transmon CCD on EY table.
  • Reverted MC autolocker to nominal, high power version.
  • Raised PSL output power back to nominal level by turning the waveplate. At the PSL shutter I measured a power of 1.03W. It occured to me too late that I realigned the PMC only afterwards and increased its transmission by a few percent, so I'll have to re-measure the actual PSL power.
  • MC is locked with its transmission back up to ~15,400 counts. The autolocker is not very good at obtaining the lock, as it seems to try to turn the VCO gain up too far and loses lock. The script probably needs a revision.
  • The YARM was pretty badly aligned. We used the green light to visually center the beam on the test masses AND had to go exploring with the TTs to see IR flashes in the first place. We got the YARM to lock to IR and were able to run the dither alignment. The maximum transmission we saw was on the order of 0.85. However, something strange is happening with the LSC control of the armlength. When the lock is engaged it drives PIT and YAW, which manifests itself in the OpLev signal and variable transmitted power on the TRY PDs. Osamu helped us diagnose this and was able to reduce the effect by tuning the POS gains to the individual ETMY OSEMS. The problem persisted even after using the new matrix diagonalization coefficients, we'll have to investigate this further and also take a look at the filters in the feedback path.
  • ITMX is still stuck and way out of alignment, so we couldn't even start with the green light in the XARM.
ELOG V3.1.3-